OpenStudy (anonymous):

if sound travels 350m/s how lon does it take to travel 110m

OpenStudy (anonymous):

350:1 110:x Therefore x = 110/350=11/35 seconds

OpenStudy (anonymous):

u mean 11.35s

OpenStudy (anonymous):

Nope. I mean 11/35 seconds. The answer is in fraction or if you want it in decimal, then it's 0.314 (3 d.p)

OpenStudy (anonymous):

thx